OpenStudy (anonymous):

8x-9y=24.5 7y-2x=-8.5 solve by elimination

OpenStudy (anonymous):

8x-9y=24.5 =8x-9y=24.5 4(-2x+7y=-8.5)=-8x+28y=-34 = 0x+19y=-9.5 y=-.5 7(-.5)-2x=-8.5 -7/2 - 2x=-8.5 -2x=-5 x=5/2 y=-.5

OpenStudy (anonymous):

how do I make it an ordered pair, a fraction ordered pair?

OpenStudy (anonymous):

-9y+8x=24.5 rearanged -9y+8x=24.5 7y-2x=-8.5 mult by 4 28y-8x=-34 add ------------- 19y =-9.5 y=-1/2 sub this to the other eq

OpenStudy (anonymous):

(x,y)=(5/2,-1/2)

OpenStudy (anonymous):

oh lol thanks you

OpenStudy (anonymous):

y.w. lol good luck now

OpenStudy (anonymous):

see i can do the first half of the problem fine, but I have a problem understanding how they come up with the fractions in the second half

OpenStudy (anonymous):

the 2nd half is substitution 7(-1/2)-2x=-8.5 -7/2 - 2x=-8.5 -2x=-5 x=5/2

OpenStudy (anonymous):

well no wonder lol. thank you so much for yor help I really appreciate it :D

OpenStudy (anonymous):

ok good luck now